Campbell, a single taxpayer, earns $400,000 in taxable in bonds. (Use the US tax rate sch and $2,000 in interest from an investment in State of New York edule) (Do not round intermediate calculations. Round Federal tax to 2 decimal places.) a. How much federal tax will she owe? b. What is her average tax rate? c. What is her effective tax rate? d. What is her current marginal tax rate? Answer is not complete. Complete this question by entering your answers in the tabs below.
media%2Ff19%2Ff19b14e5-767d-493b-9b3f-50
Answer #1
a) Federal tax owed = 46643.75+33%*(400000-191650) = 115399.25
b) Average tax rate = Total tax/Taxable income = 115399.25/400000 = 28.85%
c) Effective tax rate = Total tax/Total income = 115399.25/402000 = 28.71%
d) Current marginal tax rate is the tax rate applicable to the next $1 of taxable income, which is 33%(as income from 191650 to 416700 is taxable at 33%)